on a certain river, a paddle boat travels 8 miles upstream against the current in the same amount of time it travels 16 miles downstream with the current . If the current of the river is 2 miles per hour determine the speed of the boat in still water
Answer by TimothyLamb(4379)

s = d/t
---
b = boat speed in still water
c = current speed
c = 2 mph
---
up:
b - c = 8/t
b - 2 = 8/t
b = 8/t + 2
---
down:
b + c = 16/t
b = 16/t - 2
---
8/t + 2 = 16/t - 2
8/t = 4
t = 2 hours
---
b = 8/t + 2
b = 8/2 + 2
b = 4 + 2
---
answer:
b = 6 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php